पहली संख्या 3240,दूसरी संख्या 3600 तो तीसरी संख्या क्या होगी,जबकि तीनो का HCF 36 और LCM 4762800 है
 

Grade:10

1 Answers

Arun
5 years ago
3600=2^4*3^2*5^23240=2^2*3^4*5H.C.F =36=2^2 *3^2Since H.C.F is the product of lowest power of common Factor , so the third Number must have (2^2 *3^2) as it`s factor .Since L.C.M is the Product of Highest power of Common Prime Factor , So the Third number must have 3^5 and 7^2 as it`s factor .Therefore , Third number =2^2*3^5*7^2
 
Regards
Arun